# Servo Integrated Nonlinear Model Predictive Control for Overactuated Tiltable-Quadrotors

### v1

@@ III Control
Unlike some cascade control algorithms, our proposed NMPC controller leverages the full nonlinear dynamics as the control model, emphasizing a concise and elegant “end-to-end” style. During this section, we first describe the generation of control reference in Section III-A. Then this target is converted into a cost function inside a finite-time optimal control problem (OCP), which is the core concept of NMPC in Section III-B. One drawback of NMPC is its sensitivity to model error, which leads to steady-state error during flight. To solve this problem, a simple integral mechanism is introduced in Section III-C.
@@ III-A Generation of Control Reference
Although the NMPC controller is able to work with only ${}^{W}\bm{p}_{r}$ and ${}^{W}_{G}\bm{q}_{r}$, the optimizer can converge faster by computing reference states as comprehensively as possible. Considering that the computation cost is trivial, this calculation is worthwhile. Thus, all our subsequent experiments are conducted with the full-state reference, and the generation method is introduced here.
Compared with normal quadrotors, tiltable-quadrotors are able to independently track three-dimensional attitude at the same time. Hence, we need both reference position ${}^{W}\bm{p}_{r}$ and attitude ${}^{W}_{G}\bm{q}_{r}$ as control target. Another difference in control reference comes from the number, where NMPC requires a sequence of reference points as the control target, not just one point for traditional feedback controllers.
The control task can be divided into two types: tracking a constant point or tracking a trajectory. In our system, if the robot tracks a constant point, then all reference points are set as the same, and the reference velocity ${}^{W}\bm{v}_{r}$, acceleration ${}^{W}\dot{\bm{v}}_{r}$, angular velocity ${}^{G}\bm{\omega}_{r}$, and angular acceleration ${}^{G}\dot{\bm{\omega}_{r}}$ are set to zero. Although the research [19] claims that adding the point to the end of the reference can make optimizer easier to find a solution, we find it slows response speed. For tracking a trajectory, the reference points are calculated with a shifting prediction time interval, and ${}^{W}\bm{v}_{r}$, ${}^{G}\bm{\omega}_{r}$, ${}^{W}\dot{\bm{v}}_{r}$, and ${}^{G}\dot{\bm{\omega}_{r}}$ can be calculated by differentiating position and attitude. Then the desired ${{}^{G}\bm{f}_{u,r}}$ and ${{}^{G}\bm{\tau}_{u,r}}$ can be calculated from eqs. 9 and 11.
On the basis of these states, we can obtain the reference thrust $f_{i,r}$ and servo angle $\alpha_{i,r}$ through control allocation. The relations between reference wrench and a group of virtual input $\bm{z}$ can be expressed as
$\left[{}^{G}\bm{f}_{u,r},^{G}\bm{\tau}_{u,r}\right]^{T}=\bm{A}\ \bm{z},\quad\text{where}$ (12)
$\bm{z}=$ $\left[\begin{array}[]{c}f_{1,r,h}\\ f_{1,r,v}\\ \vdots\\ f_{N_{p},r,h}\\ f_{N_{p},r,v}\end{array}\right]=\left[\begin{array}[]{c}f_{1,r}\sin{\alpha_{1,r}}\\ f_{1,r}\cos{\alpha_{1,r}}\\ \vdots\\ f_{N_{p},r}\sin{\alpha_{N_{p},r}}\\ f_{N_{p},r}\cos{\alpha_{N_{p},r}}\end{array}\right],$ (23)
where the allocation matrix $\bm{A}$ can be derived from eqs. 5 and 6 using symbolic computation tools. Then we can calculate $f_{i,r}$ and $\alpha_{i,r}$ as follows
$\bm{z}$ $=\bm{A}^{\dagger}\ \left[{}^{G}\bm{f}_{u,r},^{G}\bm{\tau}_{u,r}\right]^{T},$ (24)
$f_{i,r}$ $=\sqrt{f_{i,r,h}^{2}+f_{i,r,v}^{2}},$ (25)
@@
where $()^{\dagger}$ means the Moore–Penrose inverse operation. Note that $\bm{A}^{\dagger}$ is unchanged and only needs to compute once.
@@ III-B Nonlinear Model Predictive Control
Nonlinear model predictive control converts the control problem to a constrained nonlinear optimization problem. For quadrotors, depending on the coordinate space, we can choose either a nonlinear model with linear constraints and a simple cost function, or a linear model with nonlinear constraints and a complex cost function. The former is easier to understand while the latter may have advantages in computation speed (please read [20] for some discussions about normal quadrotors). Considering that the former one is more intuitive and has been deployed onboard successfully in many latest research [14, 15], we select the nonlinear model version.
@@
We select the state as $\bm{x}=\left[{{}^{W}\bm{p}},{{}^{W}\bm{v}},{{}^{W}_{G}\bm{q}},{{}^{G}\bm{\omega}},{\bm{\alpha}}\right]^{T}$ and the control input as $\bm{u}=\left[\bm{f}_{c},\bm{\alpha}_{c}\right]^{T}$. Given the reference $\bm{x}_{r}$, $\bm{u}_{r}$, we define the state error as $\overline{\bm{x}}=\left[\overline{\bm{p}},\overline{\bm{v}},\mathcal{V}(\bm{q}_{e}),\overline{\bm{\omega}},\overline{\bm{\alpha}}\right]^{T}$ and the control input error as $\overline{\bm{u}}=\left[\overline{\bm{f}_{c}},\bm{\alpha}_{c}-\bm{\alpha}\right]^{T}$, where the over-line symbol denotes ()¯$\overline{(\cdot)}=(\cdot)-(\cdot)_{r}$ if no special explanation, $\mathcal{V}(\cdot)$ represents the vector part of a quaternion $\mathcal{V}(\bm{q}):=\left[q_{x},q_{y},q_{z}\right]^{T}$, and ${\bm{q}}_{e}=\bm{q}\circ\bm{q}_{r}^{-1}$ represents the quaternion error. Note that the error term for servo angle command is defined as $\overline{\bm{\alpha}}_{c}=\bm{\alpha}_{c}-\bm{\alpha}$ instead of the error w.r.t. reference $\overline{\bm{\alpha}}_{c}^{\prime}=\bm{\alpha}_{c}-\bm{\alpha}_{r}$, aiming to penalize the right-hand of eq. 7 and avoid the sudden change of servo angle. More analysis is presented in Section IV.
@@
Then the optimal control problem in NMPC can be formulated as follows
$\underset{\bm{x}_{k},\bm{u}_{k}}{\rm minimize\ }\quad$ $\sum\limits_{k=0}^{N-1}\left(\overline{\bm{x}}^{T}_{k}\bm{{Q}}\overline{\bm{x}}_{k}+\overline{\bm{u}}^{T}_{k}\bm{{R}}\overline{\bm{u}}_{k}\right)+\overline{\bm{x}}^{T}_{N}\bm{{Q}}_{N}\overline{\bm{x}}_{N},$ (27)
${\rm subject\ to}\quad$ $\bm{x}_{k+1}=f\left(\bm{x}_{k},\bm{u}_{k}\right),\quad k=0:N-1,$ (28)
$\bm{x}_{0}=\hat{\bm{x}},$ (29)
$\left\|v_{x,y,z}\right\|\leq v_{\rm limit},\quad\left\|\omega_{x,y,z}\right\|\leq\omega_{\rm limit},$ (30)
$\bm{u}_{\min}\leq\bm{u}_{k}\leq\bm{u}_{\max},$ (31)
@@
where $\bm{Q}$, $\bm{R}$, $\bm{Q}_{N}$ are positive diagonal matrices representing weights for state cost, control energy cost, and terminal cost, respectively. The eq. 28 indicates the dynamics constraint, where $f(\cdot)$ refers to the full tiltable-quadrotor model established from eq. (2-11). This model is discretized by the fourth-order Runge–Kutta method with $t_{\rm integ}$ integrating time. The eq. 29 denotes the initial value constraint, a critical component for feedback in NMPC, wherein $\hat{\bm{x}}$ represents the estimated state obtained from the estimator. Finally, the eqs. 30 and 31 refer to the state and input constraints, respectively, which is mainly decided from physical limits and safety concern.
Equation 27 is an nonlinear least-square cost function, where the nonlinearity comes from quaternions. Recall the quaternion error is denoted as ${\bm{q}}_{e}=\bm{q}\circ\bm{q}_{r}^{-1}$, then its cost term can be written as
$\overline{\bm{q}}^{T}_{k}\bm{Q}_{q}\overline{\bm{q}}_{k}=\left\|{\rm sgn}(q_{e,w})\cdot\mathcal{V}({\bm{q}}_{e})\right\|_{\bm{Q}}^{2}=\mathcal{V}({\bm{q}}_{e})^{T}\bm{Q}_{q}\mathcal{V}({\bm{q}}_{e}).$ (32)
When NMPC is executing, the finite-time OCP problem defined above is solved during each control loop in $t_{s}$ frequency, and traditionally the first control command is extracted from the optimal sequence as the control input. However, this way neglects the time caused by NMPC computation and other delay factors, making the command out-of-date. To address this challenge, recognizing the nonlinearity is small within short time intervals, we predict the control command forward using linear interpolation as follows
$\bm{u}_{\rm now}=\frac{t_{s}+t_{d}}{t_{\rm integ}}(\bm{u}^{*}_{1}-\bm{u}^{*}_{0})+\bm{u}^{*}_{0},$ (33)
where $\bm{u}^{*}_{0}$ and $\bm{u}^{*}_{1}$ are the first two commands from the generated control sequence, $t_{s}$ is the time for control loop, $t_{d}$ is the dead time of actuators, and $t_{\rm integ}$ is the time for one prediction step. At the beginning of the next control loop, the final input $\bm{u}_{\rm now}$ is transmitted to low-level autopilot for execution.
@@
In implementation, we leverage several common techniques in NMPC community to accelerate computation. Specifically, warm-starting, real-time iteration (RTI), and multiple-shooting are adopted. The warm-starting tries to accelerate by giving an initial guess near the final solution, where the guess comes from the last round’s solution. The RTI only calculates the sequential quadratic programming (SQP) once in one control iteration, preferring speed instead of optimality. Finally, the multiple-shooting divides the original OCP into several smaller optimization problems and try to solve them in parallel. We recommend [19] to interested readers for more details.
III-C Integral Term for Model Error
To alleviate the influence of model error to NMPC, we adopt an integral term as follows to compensate for pose error
${{}^{W}\bm{f}_{d}}$ $=\text{ITerm}\left({{}^{W}\hat{\bm{p}}}-{{}^{W}\bm{p}_{r}}\right),$ (34)
${{}^{G}\bm{\tau}_{d}}$ $=\text{ITerm}\left(\text{sgn}(q_{e,w})\cdot\mathcal{V}(\hat{\bm{q}}\circ\bm{q}_{r}^{-1})\right),$ (35)
@@ III-C Integral Term for Model Error
where $\text{sgn}(\cdot)$ denotes sign function. The digital implementation of integral term $u[k+1]=\text{ITerm}(e[k+1])$ with trapezoidal rule and anti-windup is given as [21]
@@
The compensating wrench ${{}^{W}\bm{f}_{d}}$, ${{}^{G}\bm{\tau}_{d}}$ is calculated before each NMPC running and then transmitted as inputs into the optimization process through eqs. 9 and 11.
@@ IV Simulation
In this section, we present simulation results to reveal the impact of different NMPC designs on the final performance. To eliminate the influence of model error, disturbance, and other unknown effects, we implement an ideal simulation environment without model error and noise. In the subsequent comparisons, we choose the control frequency as 100Hz to align with real-world experiments, and we choose the simulation frequency as 200Hz since our estimator for real robot is running at this frequency. The simulating model is always the full nonlinear model with servo dynamics eq. (2-11). The weights are all set as Table I if no special explanation.
IV-A The Effect of Servo Dynamics to Convergence
Aiming to explore the effect of servo dynamics eq. 7 to NMPC, we implement two kinds of NMPC: one has no servo dynamics and hence has no servo angle in states, as well as the other one includes servo dynamics. Two NMPC controllers are commanded to takeoff and reach $1$m in height. The result is shown in Fig. 4.
(a) NMPC without servo model, $R_{\alpha}=1$
(b) NMPC without servo model, $R_{\alpha}=5$
(c) NMPC with servo model, $t_{\rm servo}=0.0859{\rm s}$, $Q_{\alpha}=R_{\alpha}=5$
(d) NMPC with servo model, $t_{\rm servo}=0.045{\rm s}$, $Q_{\alpha}=R_{\alpha}=5$
Figure 4: Comparative analysis of NMPC performance with and without the servo model during a takeoff mission. The actual $t_{\rm servo}$ in simulation is always 0.0859s. Figure (a) and (b) demonstrate that increasing the cost for servo commands mitigates oscillation. Figure (b) and (c) illustrate that incorporating the servo model results in smoother controller performance. However, as shown in (d), an overly small $t_{\rm servo}$ can hinder the NMPC’s convergence.
@@ IV-A The Effect of Servo Dynamics to Convergence
It is natural to think that adding a model increases the OCP complexity, resulting higher difficulty in convergence. However, as revealed by Fig. 4, the NMPC with servo model achieves less oscillation and faster convergence. We believe this is due to the narrower search space implicitly introduced by the first-order servo model. Without this model, the next servo angle command can be any value within the physical limits. With this model, the command is limited in the nearby area of the real servo angle, and this extra limitation accelerates the convergence. For the accuracy of time constant, too short value reduces the narrowing effect and leads to oscillation in Fig. 4d, while the correct or larger values achieves the similar response shown in Fig. 4c. Although the no-servo version achieves convergence in the end, we report that it diverges and crashes in noise-existed Gazebo simulation, not mention in real world. Hence, the inclusion of servo model is not only a plus for control performance, but also a must to achieve real-world usage.
IV-B The Effect of Servo Dynamics Cost
We want to compare the result of traditional cost $\overline{\bm{\alpha}}_{c}^{\prime}=\bm{\alpha}_{c}-\bm{\alpha}_{r}$ and our proposed cost $\overline{\bm{\alpha}}_{c}=\bm{\alpha}_{c}-\bm{\alpha}$ under the influence of inaccurate $\bm{\alpha}_{r}$. This inaccuracy happens in reality due to the model error or the absent of allocation module. We set the reference $\bm{\alpha}_{r}=\bm{0}$, set $Q_{\alpha}=1$, and try to use NMPC to track the attitude change. The result is presented in Fig. 5.
(a) Old $\overline{\bm{\alpha}}_{c}^{\prime}=\bm{\alpha}_{c}-\bm{\alpha}_{r}$, $R_{\alpha}=5$
(b) Old $\overline{\bm{\alpha}}_{c}^{\prime}=\bm{\alpha}_{c}-\bm{\alpha}_{r}$, $R_{\alpha}=20$
(c) New $\overline{\bm{\alpha}}_{c}=\bm{\alpha}_{c}-\bm{\alpha}$, $R_{\alpha}=5$
(d) New $\overline{\bm{\alpha}}_{c}=\bm{\alpha}_{c}-\bm{\alpha}$, $R_{\alpha}=20$
Figure 5: Comparison of orientation tracking performance with inaccurate $\bm{\alpha}_{r}$. Compared with the old cost function, the new cost effectively mitigates the adverse effect of inaccuracies in $\bm{\alpha}_{r}$.
From the figure, the tracking error of old cost is sensitive to $R_{\alpha}$, while the new cost always maintains a small error. The traditional setting ($\overline{\bm{\alpha}}_{c}^{\prime}=\bm{\alpha}_{c}-\bm{\alpha}_{r}$ and $\overline{\bm{\alpha}}=\bm{\alpha}-\bm{\alpha}_{r}$) mixes two missions in $\overline{\bm{\alpha}}_{c}^{\prime}$: tracking $\bm{\alpha}_{r}$ and indirectly reducing the difference with actual angle $\bm{\alpha}$. On the contrary, the new setting ($\overline{\bm{\alpha}}_{c}=\bm{\alpha}_{c}-\bm{\alpha}$ and $\overline{\bm{\alpha}}=\bm{\alpha}-\bm{\alpha}_{r}$) makes $\overline{\bm{\alpha}}_{c}$ only control the difference with $\bm{\alpha}$, and purely leaves the tracking task to $\overline{\bm{\alpha}}$. If the reference $\bm{\alpha}$ is accurate, two settings achieve the same performance. Yet if $\bm{\alpha}$ is inaccurate, this functional decomposition ensures that a large $R_{\alpha}$ does not cause large steady error. This is quite supportive when $R_{\alpha}$ is set to a large number to reduce servo oscillation in real world.

### v2

@@ III Control
Unlike some cascade control algorithms, our proposed NMPC controller leverages the fully nonlinear dynamics as the control model, emphasizing a concise and elegant “end-to-end” style. During this section, we first describe the generation of control reference in Sec. III-A. Then this target is converted into a cost function inside a finite-time optimal control problem (OCP), which is the core concept of NMPC in Sec. III-B. One drawback of NMPC is its sensitivity to model error, which leads to a steady-state error during flight, especially influenced by the ground effect in the Z axis. To solve this problem, a simple integral unit is introduced in Sec. III-C.
@@ III-A Generation of Control Reference
Although NMPC can work with only ${}^{W}\bm{p}_{r}$ and ${}^{W}_{G}\bm{q}_{r}$, the optimizer can converge faster by computing reference states as comprehensively as possible. This calculation is worthwhile since the computation cost is trivial. Thus, all our subsequent experiments are conducted with the full-state reference, and the generation method is introduced here.
Compared with normal quadrotors, tiltable-quadrotors are able to independently track three-dimensional attitudes at the same time. Hence, we need both reference position ${}^{W}\bm{p}_{r}$ and attitude ${}^{W}_{G}\bm{q}_{r}$ as control targets. Another difference in control reference comes from the number, where NMPC requires a sequence of reference points as the control target, not just one point for traditional feedback controllers.
The control task can be divided into two types: tracking a constant point or tracking a trajectory. In our system, if the robot tracks a constant point at time $t$, then all reference points after $t$ are set the same, and the reference velocity ${}^{W}\bm{v}_{r}$, acceleration ${}^{W}\dot{\bm{v}}_{r}$, angular velocity ${}^{G}\bm{\omega}_{r}$, and angular acceleration ${}^{G}\dot{\bm{\omega}_{r}}$ are set to zero. For tracking a trajectory, the reference points are calculated with a shifting prediction time interval, and ${}^{W}\bm{v}_{r}$, ${}^{W}\dot{\bm{v}}_{r}$, ${}^{G}\bm{\omega}_{r}$, and ${}^{G}\dot{\bm{\omega}_{r}}$ can be calculated by differentiating position and attitude. Then the desired ${{}^{G}\bm{f}_{u,r}}$ and ${{}^{G}\bm{\tau}_{u,r}}$ can be calculated from (6b) and (6d), where the gyroscopic term is omitted for simplicity.
On the basis of these states, we can obtain the reference thrust $f_{i,r}$ and servo angle $\alpha_{i,r}$ through control allocation. The relations between the reference wrench and a group of virtual input $\bm{z}$ can be expressed as
$\left[{}^{G}\bm{f}_{u,r},^{G}\bm{\tau}_{u,r}\right]^{T}=\bm{A}\ \bm{z},\quad\text{where}$ (7a)
$\bm{z}=$ $\left[\begin{array}[]{c}f_{1,r,h}\\ f_{1,r,v}\\ \vdots\\ f_{N_{p},r,h}\\ f_{N_{p},r,v}\end{array}\right]=\left[\begin{array}[]{c}f_{1,r}\sin{\alpha_{1,r}}\\ f_{1,r}\cos{\alpha_{1,r}}\\ \vdots\\ f_{N_{p},r}\sin{\alpha_{N_{p},r}}\\ f_{N_{p},r}\cos{\alpha_{N_{p},r}}\end{array}\right],$ (7l)
where the allocation matrix $\bm{A}$ can be derived from (4a) and (4b) using symbolic computation tools. Then we can calculate $f_{i,r}$ and $\alpha_{i,r}$ as follows
$\bm{z}$ $=\bm{A}^{\dagger}\ \left[{}^{G}\bm{f}_{u,r},^{G}\bm{\tau}_{u,r}\right]^{T},$ (8a)
$f_{i,r}$ $=\sqrt{f_{i,r,h}^{2}+f_{i,r,v}^{2}},$ (8b)
@@
where $()^{\dagger}$ means the Moore–Penrose inverse operation. Note that $\bm{A}^{\dagger}$ is unchanged and only needs to be computed once.
As revealed by [13], NMPC has superiority in tracking dynamically infeasible trajectories. Hence, our requirement in this part is a middle-quality reference with a trivial computational burden, which has the potential for online replanning. The pose reference can be generated from parametric equations. If several pose points are given, the reference can be also obtained from the minimum-acceleration planning method [20] or other evolutionary versions.
@@ III-B Nonlinear Model Predictive Control
Nonlinear model predictive control converts the control problem into a constrained nonlinear optimization problem. For quadrotors, depending on the coordinate space, we can choose either a nonlinear model with linear constraints and a simple cost function, or a linear model with nonlinear constraints and a complex cost function. The former is easier to understand while the latter may have advantages in computational speed (please read [21] for discussions about normal quadrotors). Considering that the former one is more intuitive and has been deployed onboard successfully in many recent research [13, 14], we select the nonlinear model version.
@@
We select the state as $\bm{x}=\left[{{}^{W}\bm{p}},{{}^{W}\bm{v}},{{}^{W}_{G}\bm{q}},{{}^{G}\bm{\omega}},{\bm{\alpha}}\right]^{T}$ and the control input as $\bm{u}=\left[\bm{f}_{c},\bm{\alpha}_{c}\right]^{T}$. Given the reference $\bm{x}_{r}$, $\bm{u}_{r}$, we define the state error as $\overline{\bm{x}}=\left[\overline{\bm{p}},\overline{\bm{v}},\mathcal{V}(\bm{q}_{e}),\overline{\bm{\omega}},\overline{\bm{\alpha}}\right]^{T}$ and the control input error as $\overline{\bm{u}}=\left[\overline{\bm{f}_{c}},\bm{\alpha}_{c}-\bm{\alpha}\right]^{T}$, where the over-line symbol denotes ()¯$\overline{(\cdot)}=(\cdot)-(\cdot)_{r}$ if no special explanation, $\mathcal{V}(\cdot)$ represents the vector part of a quaternion $\mathcal{V}(\bm{q}):=\left[q_{x},q_{y},q_{z}\right]^{T}$, and ${\bm{q}}_{e}=\bm{q}\circ\bm{q}_{r}^{-1}$ represents the quaternion error. Note that the error term for the servo angle command is defined as $\overline{\bm{\alpha}}_{c}=\bm{\alpha}_{c}-\bm{\alpha}$ instead of the error w.r.t. reference $\overline{\bm{\alpha}}_{c}^{\prime}=\bm{\alpha}_{c}-\bm{\alpha}_{r}$, aiming to penalize the right-hand of (5) to avoid the sudden change of servo commands.
@@
Then the optimal control problem in NMPC can be formulated as a nonlinear least-square problem
$\underset{\bm{x}_{k},\bm{u}_{k}}{\rm minimize\ }\quad$ $\sum\limits_{k=0}^{N-1}\left(\overline{\bm{x}}^{T}_{k}\bm{{Q}}\overline{\bm{x}}_{k}+\overline{\bm{u}}^{T}_{k}\bm{{R}}\overline{\bm{u}}_{k}\right)+\overline{\bm{x}}^{T}_{N}\bm{{Q}}_{N}\overline{\bm{x}}_{N},$ (9a)
${\rm subject\ to}\quad$ $\bm{x}_{k+1}=\bm{f}\left(\bm{x}_{k},\bm{u}_{k}\right),\quad k=0:N-1,$ (9b)
$\bm{x}_{0}=\hat{\bm{x}},$ (9c)
$\left\|v_{x,y,z}\right\|\leq v_{\rm limit},\quad\left\|\omega_{x,y,z}\right\|\leq\omega_{\rm limit},$ (9d)
$\bm{u}_{\min}\leq\bm{u}_{k}\leq\bm{u}_{\max},$ (9e)
@@
where $\bm{Q}$, $\bm{R}$, $\bm{Q}_{N}$ are positive diagonal matrices representing weights for state cost, control energy cost, and terminal cost, respectively. The (9b) indicates the dynamics constraint, where $\bm{f}(\cdot)$ refers to the full tiltable-quadrotor model established from (2) to (6). This model is discretized by the fourth-order Runge–Kutta method with $t_{\rm integ}$ integrating time. The (9c) denotes the initial value constraint, a critical component for feedback in NMPC, wherein $\hat{\bm{x}}$ represents the estimated state from the estimator. Finally, the (9d) and (9e) refer to the state and input constraints, respectively, which are mainly decided from safety concerns and physical limits.
After calculation, the first element of the optimized sequence $\bm{u}^{*}$ is transmitted to a low-level autopilot for execution:
${\color[rgb]{0,0,0}\bm{u}_{\rm now}=\bm{u}^{*}_{0}.}$ (10)
@@
In implementation, we leverage several typical techniques in the NMPC community to accelerate computation. Specifically, warm-starting, real-time iteration (RTI), and multiple-shooting are adopted. The warm-starting tries to accelerate by giving an initial guess near the final solution, where the guess comes from the last round’s result. The RTI only calculates the sequential quadratic programming (SQP) once in one control iteration, preferring speed instead of optimality. Finally, the multiple-shooting divides the original OCP into several smaller optimization problems and tries to solve them in parallel. We recommend [22] to interested readers for more details.
III-C Integral Term for Ground Effect
When rotor-based drones fly near the ground, the aerodynamic interaction between the drone’s propellers and the environment can increase the lift, resulting in higher force in the Z axis. We empirically find that the model error caused by this phenomenon is much larger than other axes, so we adopt an integral term as follows to compensate for the Z error
${\color[rgb]{0,0,0}f_{d,z}=\text{ITerm}\left({{}^{W}\hat{p}_{z}}-{{}^{W}p_{z,r}}\right).}$ (11)
@@ III-C Integral Term for Ground Effect
The digital version of the integral term $u[k+1]=\text{ITerm}(e[k+1])$ with trapezoidal rule and anti-windup is given as [23]
@@
The compensating force ${{}^{W}\bm{f}_{d}}=\left[0,0,f_{d,z}\right]^{T}$ is calculated before each NMPC round and then transmitted as parameters into the optimization process through (6b).
(a) Pose tracking for NMPC without servo and thrust model
(b) Pose tracking for NMPC with servo and without thrust model.
(c) Pose tracking for NMPC with servo and thrust model
Figure 4: Comparative analysis of NMPC considering different models in an ideal simulation, where the control targets are described in the main text. The drastic command oscillation in (a) leads to optimization errors in noisy simulations such as Gazebo, and incorporating a servo model in (b) significantly mitigates this phenomenon. From (b) and (c), adding a thrust model merely results in a smoother tracking performance. The result of only considering the thrust is similar to (a).
@@ IV Simulation
In this section, we perform simulations to reveal the impact of the servo model on the NMPC performance. To eliminate the influence of model error, disturbance, and other unknown effects, we implement an ideal simulation environment without model error and noise.
[…]
IV-A The Influence of Servo Model on Optimization Convergence
@@
It is natural to think that adding a model increases the OCP complexity, resulting in higher difficulty in convergence. However, as revealed by Fig. 4a and Fig. 4b, the servo-integrated NMPC achieves less oscillation and faster optimization convergence. We believe this is due to the narrower search space introduced implicitly by the first-order servo model. Without this model, the next servo angle command can be any value within the physical limits. With this model, the command is limited in the nearby area of the real servo angle, and this extra limitation accelerates the convergence. Although the no-servo version achieves optimization convergence in the end, we report it diverging and crashing in noise-existing simulation, not mentioned in the real world. Hence, the inclusion of the servo model is a must for real-world usage.
We also observe that increasing the servo angle’s weight to a large number (above 50) can eliminate the oscillation. However, this way relies heavily on the accuracy of $\bm{\alpha}_{r}$ and leads to unacceptable error when tracking infeasible trajectories.
IV-B The Influence of Servo Model on Flight Performance
On the basis of optimization convergence, considering the servo dynamics can reduce the fluctuations and hence increase the flight performance as revealed by Fig. 4a and Fig. 4b.
[…]
Although considering the thrust model can theoretically improve performance, focusing solely on the servo model is easier for implementation. Adding the thrust model to NMPC needs more computation, and also requires extra state-feedback of thrust, which can be calculated from rotor speed through (1). Many servos support angle measurement, but few ESCs can measure rotor speed at high frequencies. Therefore, we recommend only considering the servo model at the beginning. If oscillations are observed as in Fig. 4b, modeling the thrust should then be considered.